Two domestic parcel companies collect their charges as follows: Company A collects $5 for every pound and a service charge of $3. Company B collects $7 for every pound and a service charge of $2. For what weight of the parcel is the charge same in both companies?

Answer by stanbon(75887)

Company A collects $5 for every pound and a service charge of $3. Company B collects $7 for every pound and a service charge of $2. For what weight of the parcel is the charge same in both companies?
-----------
C(A) = 3 + 5x
C(B) = 2 + 7x
-------
Equation:
3+5x = 2+7x
2x = 1
x = 1/2 lb
--------------------
Cheers,
Stan H.
===============